OpenStudy (halfdeafdarling):

Okay so just seeing if i have this right, solving for s=4v^2 and I have v=sqrts/4

OpenStudy (halfdeafdarling):

\[v=\sqrt{S/5}\]

zepdrix (zepdrix):

Solving for v? Start by isolating v^2. v^2 is being `multiplied` by 4, so we'll undo this operation by dividing by 4,\[\large\rm \frac{s}{4}=v^2\]Then apply square root to undo the `square` on v,\[\large\rm \sqrt{\frac s4}=\pm v\]When you take the square root of a square, you get two results, a positive and negative case. If this is for some type of physics class though, where v can only be a positive quantity, then you can ignore the plus/minus symbol in front. Otherwise, move the plus/minus to the other side,\[\large\rm \pm\sqrt{\frac s4}=v\]

OpenStudy (halfdeafdarling):

okay i see just double checking thank you so much